OpenStudy (anonymous):

A box has 5 beads of the same size, but all are different colors. Tina draws a bead randomly from the box, notes its color, and then puts the bead back in the box. She repeats this 3 times. What is the probability that Tina would pick a red bead on the first draw, then a green bead, and finally a red bead again?

OpenStudy (anonymous):

@deadpoolroxz @iGreen

OpenStudy (anonymous):

@deadpoolroxz

OpenStudy (anonymous):

u there?

OpenStudy (anonymous):

@deadpoolroxz

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

ill join the deadpool club but how

OpenStudy (anonymous):

lol then ill mention u in every question

OpenStudy (igreen):

Okay, 5 beads, 5 different colors.

OpenStudy (igreen):

What's the probability of picking a red?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1 of 5?

OpenStudy (anonymous):

is it righT?

OpenStudy (igreen):

Correct, what about picking a green?

OpenStudy (anonymous):

1 of 5

OpenStudy (igreen):

Correct, and another red is 1/5 \(\dfrac{1}{5} \times \dfrac{1}{5} \times \dfrac{1}{5} \)

OpenStudy (anonymous):

right

OpenStudy (igreen):

Multiply. \(\dfrac{1}{5 \times 5 \times 5}\)

OpenStudy (igreen):

5 * 5 * 5 = ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

well u take a while to type

OpenStudy (anonymous):

cant blam

OpenStudy (anonymous):

lol anyone give me an answer

OpenStudy (anonymous):

thank you dead pool heex for the medal

OpenStudy (anonymous):

you wecome you earnd it

OpenStudy (igreen):

That's wrong.. @pandaluvs

OpenStudy (igreen):

5 * 5 * 5 isn't 15..

OpenStudy (anonymous):

@iGreen sorry i said it wrong

OpenStudy (anonymous):

125

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

thats it

OpenStudy (igreen):

Yes, so our answer is \(\dfrac{1}{125}\)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

if i new how i would give you anthor medel

OpenStudy (anonymous):

http://akk.li/pics/anne.jpg if you want to see something funny hit the link

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

i saw it and im not falling for it the jeff the killer